OpenStudy (matt6288):

A number is divided by 3. Then, 5 is added to the quotient, and the result is 8. What is the number?

OpenStudy (unklerhaukus):

\[\frac n3+5=8\]

OpenStudy (anonymous):

n/3 = 3 n = 9